x power 2 + y power 2=6xy then find 2 log (x-y)

Answer by malglu(63)

x^2 + y^2 = (x-y) (x-y) +2xy
so rewrite the equation as
(x-y)^2 +2xy = 6xy
(x-y)^2 - 4xy
take logs
log (x-y)^2 = log 4xy
2 log (x-y) = log 4 + log x +log y
which is your answer